OpenStudy (anonymous):

Freddy brought some candy and gave 1/3 to Liza. Liza gave the rest to Maria. If Liza got 18 How much did Maria get?

OpenStudy (anonymous):

wouldnt you just take 18 divided by 1/3? and you have your answer?

OpenStudy (anonymous):

liza gave the rest or freddy??

OpenStudy (anonymous):

If 18 is 1/3, then multiply by three to get the total (64) and then subtract 18 to get what was given to maria (64-18 = 36)

OpenStudy (anonymous):

No to Maria.

OpenStudy (anonymous):

Freddy gave 1/3 to lisa, and he gave the rest to maria. That means he gave 2/3 to Maria. Maria is going to get twice as much as Lisa (because 2*(1/3) is 2/3) So 18*2 = 36

OpenStudy (anonymous):

as you can see there are a couple of ways to do this.

hero (hero):

Isn't 54 - 18 = 36 ?

OpenStudy (anonymous):

Thanks everyone.

hero (hero):

How'd the other guy get 36?

OpenStudy (anonymous):

But how did u get 54?

hero (hero):

18*3 = 54

OpenStudy (anonymous):

he mutliplied wrong, he should have gotten 54. 3*18 is 54

OpenStudy (anonymous):

Ouch, embarrassing. Sorry about that. With the correct multiplication that works. :(

OpenStudy (anonymous):

That okay people make mistakes.

OpenStudy (anonymous):

Find the missing # to n, 78,65,82,91,63,89,78, n. Mean= 80

OpenStudy (anonymous):

256?

hero (hero):

640, my bad

OpenStudy (anonymous):

Is that the answer?

hero (hero):

no, that's what it has to add up to

OpenStudy (anonymous):

94?

hero (hero):

Yes

OpenStudy (anonymous):

Thanks.